### v1

# Normalization of Puiseux hypersurfaces

Abstract.
It is known that the normalization of a quasi-ordinary complex singularity is a Hirzebruch-Jung, see [perez2000singularites, popescu2004analytical, aroca2005normal]. We extend this result to Puiseux hypersurfaces. Moreover, we prove that Hirzebruch-Jung singularities are precisely normalizations of Puiseux hypersurfaces. Our result holds over an algebraically closed field whose characteristic does not divide the degree of the polynomial defining the hypersurface. Finally, in the analytic complex case, we conclude that the normalization of an irreducible Puiseux hypersurface is the normalization of a complex analytic quasi-ordinary singularity.
Key words and phrases:
Puiseux hypersurfaces, quasi-ordinary singularities, toric varieties, Hirzebruch-Jung singularities
This work has been partially supported by Laboratorio Internacional Solomon Lefschetz, by DGAPA- Universidad Nacional Autónoma de México through PAPIIT IN113323, and by SECIHTI, Mexico CVU’s: 1337652, 781791, 1044169.
We are grateful to M. Spivakovsky for his helpful comments and suggestions, which greatly improved this work. The corresponding author also thanks P. Popescu-Pampu and P. González-Pérez for providing valuable references.
Corresponding Author. E-mail: damian.ochoa@im.unam.mx
Introduction
We denote by $\mathbb{K}$ an algebraically closed field. The ring of Puiseux series is defined as
$\mathbb{K}[[X^{*}]]=\mathbb{K}[[x_{1}^{*},\ldots,x_{n}^{*}]]:=\bigcup_{m\in\mathbb{Z}_{>0}}\mathbb{K}\left[\!\left[x_{1}^{1/m},\ldots,x_{n}^{1/m}\right]\!\right].$
A monic polynomial $f\in\mathbb{K}[[X]][y]$ is called a Puiseux polynomial if there exist $\xi_{1},\ldots,\xi_{d}\in\mathbb{K}[[X^{*}]]$ such that
$f=\prod_{i=1}^{d}(y-\xi_{i}).$
We consider Puiseux hypersurfaces $f$ such that $\operatorname{char}(\mathbb{K})$ does not divide $\text{deg}\,(f)$.
An analytic germ of dimension $n$ is quasi-ordinary when it is a local covering of $\mathbb{K}^{n}$ unramified outside the coordinate hyperplanes. In the complex analytic case, the Abhyankar–Jung theorem [abhyankar1955ramification] states that quasi-ordinary hypersurfaces are Puiseux. Quasi-ordinary singularities have been widely studied; see, for example, [alonso1989algorithm, MR2957197, mourtada2015polyhedral]. For an introductory text, we refer to [da2011semigrupo]. Hironaka defined a more general class of singularities containing the quasi-ordinary ones, which he called $\nu$–quasi-ordinary [hironaka2003theory]; see also [cassou2011nu, aroca2024nu].
It is natural to ask which properties can be extended from quasi-ordinary to Puiseux hypersurfaces. For instance, Lipman and Gau [lipman1988topological] associated to a root of a quasi-ordinary polynomial an ordered set of $n$-tuples of nonnegative rational numbers, called characteristic exponents. Let $\lambda^{(1)},\ldots,\lambda^{(g)}$ be the set of characteristic exponents of a quasi-ordinary root $\xi$. Then
(1) $\mathbb{K}((X))(\xi)=\mathbb{K}((X))(X^{\lambda^{(1)}},\ldots,X^{\lambda^{(g)}}).$
In [tornero2008kummer], Tornero proved that if $\xi$ is a root of a Puiseux polynomial, there exist $\lambda^{(1)},\ldots,\lambda^{(g)}$ such that (1) holds. Such a set is called a set of distinguished exponents of $\xi$. In the quasi-ordinary case, it is known that the characteristic exponents determine the topological type of the singularity. The relationship between the topological type of a Puiseux hypersurface and its distinguished exponents is not yet known.
One of the most important properties of a quasi-ordinary singularity is that its normalization is Hirzebruch–Jung, that is, analytically isomorphic to the zero-dimensional orbit of an affine toric variety defined by a maximal simplicial pair [perez2000singularites, aroca2005normal, popescu2004analytical]. This paper is devoted to showing that this property also holds for Puiseux hypersurfaces (see Theorem 5.2).
The paper is organized into five sections. In Sections 14, we review the basic notions of affine semigroups and their saturations, introduce the ring of series supported on a pointed affine semigroup, recall key results on toric varieties with precise references, and present Puiseux series together with a method to compute their distinguished exponents. Finally, in Section 5, we prove (see Lemma 5.1) that for $\xi\in\mathbb{K}[[X^{*}]]$ with distinguished exponents $\lambda^{(1)},\ldots,\lambda^{(g)}$, the normalization of $\mathbb{K}[[X]][\xi]$ coincides with the normalization of $\mathbb{K}[[X]][X^{\lambda^{(1)}},\ldots,X^{\lambda^{(g)}}]$, which allows us to establish:
Theorem [5.2] The normalization of an irreducible Puiseux hypersurface is either nonsingular or has a Hirzebruch–Jung singularity. Conversely, every Hirzebruch–Jung singularity is analytically isomorphic to the normalization of a Puiseux hypersurface.
As a corollary, using a result of Popescu-Pampu [popescu2003higher], in the complex analytic case we conclude that the normalization of an irreducible Puiseux hypersurface coincides with the normalization of a complex analytic quasi-ordinary singularity (see Corollary 5.6).
1. Affine Semigroups
A semigroup is a nonempty set equipped with an associative and commutative binary operation. A monoid is a semigroup with an identity element. A monoid is said to be pointed when its only invertible element is the identity.
Let $S$ be a monoid. Given $E\subseteq S$, the group generated by $E$ is denoted by
$\mathbb{Z}E:=\left\{\sum_{e\in\Lambda\subseteq E}c_{e}e\ \Big|\ c_{e}\in\mathbb{Z},\ \#\Lambda<\infty\right\},$
and the monoid generated by $E$ by
$\mathbb{N}E:=\left\{\sum_{e\in\Lambda\subseteq E}c_{e}e\ \Big|\ c_{e}\in\mathbb{Z}_{\geq 0},\ \#\Lambda<\infty\right\}\subseteq S\cup\{0\}.$
A monoid $S$ is said to be finitely generated if there exists a finite set $E$ such that $S=\mathbb{N}E$.
A lattice is a group isomorphic to $\mathbb{Z}^{r}$. An affine semigroup is a finitely generated monoid that can be embedded in a lattice.
A subset $\sigma$ of the form
$\sigma=\langle E\rangle:=\left\{\sum_{e\in\Lambda\subseteq E}c_{e}e\ \Big|\ c_{e}\in\mathbb{R}_{\geq 0},\ \#\Lambda<\infty\right\}$
is called a cone. The set $E$ is said to be a system of generators of $\sigma$. When a cone has a finite system of generators $E\subseteq\mathbb{Z}^{n}$, it is called a rational convex polyhedral cone. A cone is strongly convex when it does not contain a one-dimensional linear subspace.
Lemma 1.1 (Gordan).
Let $\sigma\subseteq\mathbb{R}^{n}$ be a rational convex polyhedral cone. Then $S=\sigma\cap\mathbb{Z}^{n}$ is a finitely generated semigroup.
See [cox2024toric, Proposition 1.2.17] and [Kempf, page 7].
Definition 1.2.
A monoid $S$ is said to be saturated if for all $k\in\mathbb{N}$ and $m\in\mathbb{Z}S$, the condition $km\in S$ implies that $m\in S$. The saturation $\bar{S}$ of a monoid $S$ is the smallest saturated monoid containing $S$.
Example 1.3.
The monoid $S$ generated by $\{2,3\}$ is not saturated. Indeed, taking $k=2$ and $m=1$, we have $km=2\in S$ while $m\notin S$. The saturation of $S$ is $\mathbb{Z}_{\geq 0}$.
Lemma 1.4.
Let $S$ be a semigroup of $\mathbb{Z}^{n}$, let $M=\mathbb{Z}S$, and let $\sigma$ be the cone generated by $S$ in $\mathbb{R}^{n}$. Then
$\bar{S}=M\cap\sigma.$
See [cox2024toric, Proposition 1.3.8] and [Kempf, Theorem 1’, Chapter 1, Section 1].
We conclude this section by stating some equivalences between semigroups and cones.
Remark 1.5.
Given a monoid $S\subseteq M$, with $M\simeq\mathbb{Z}^{n}$, we have:
  1. (1)
    $S$ is affine if and only if $\langle S\rangle$ is rational.
  2. (2)
    $S$ is pointed and finitely generated if and only if $\langle S\rangle$ is rational and strongly convex.
  3. (3)
    $S$ is saturated if and only if $S=\langle S\rangle\cap\mathbb{Z}S.$
2. Normal Toric Varieties
An affine toric variety is an irreducible affine variety $V$ containing a torus $T\simeq(\mathbb{K}^{*})^{n}$ as a Zariski open subset, such that the algebraic action of $T$ on itself extends to an action of $T$ on $V$.
Given an affine semigroup $S$, we denote by $\mathbb{K}[S]$ the semigroup algebra
$\mathbb{K}[S]:=\Big\{\sum_{\alpha\in\Lambda}a_{\alpha}\mathbf{x}^{\alpha}\ \Big|\ a_{\alpha}\in\mathbb{K},\ \Lambda\subseteq S,\ \#\Lambda<\infty\Big\}.$
Proposition 2.1.
$S$ is an affine semigroup if and only if $\mathbb{K}[S]$ is the coordinate ring over $\mathbb{K}$ of an affine toric variety.
See [cox2024toric, Propositions 1.1.14 and 1.3.5], [Kempf, Theorem 1, Chapter 1, Section 1].
By Proposition 2.1 and Lemma 1.1, given a rational convex polyhedral cone $\sigma\subseteq\mathbb{R}^{n}$ and a lattice $M\simeq\mathbb{Z}^{n}$, one can construct an affine toric variety
$T_{\sigma}=\mathrm{Spec}\,\mathbb{K}[\sigma\cap M].$
In the following, we consider $\sigma$ to be a strongly convex rational cone of maximal dimension in $\mathbb{R}^{n}$.
Definition 2.2.
An affine variety $\mathbb{X}$ is said to be normal if its coordinate ring $\mathcal{A}(\mathbb{X})$ is integrally closed in its field of fractions. The normalization $\mathcal{N}(I)$ of a domain $I$ is the integral closure of $I$ in its field of fractions.
Example 2.3.
The toric variety $\mathbb{X}=V(x^{3}-y^{2})$ is not normal: The monic polynomial
$p(z)=z^{2}-\bar{x}\in\mathcal{A}(\mathbb{X})[z].$
has as root $\alpha=\frac{\bar{y}}{\bar{x}}$ that is in the field of fractions of $\mathcal{A}(\mathbb{X})$, but it does not lie in $\mathcal{A}(\mathbb{X})$.
Lemma 2.4.
Let $\bar{S}$ be the saturation of the affine semigroup $S$. The normalization of $\mathbb{K}[S]$ is $\mathbb{K}[\bar{S}]$.
See  [cox2024toric, Propositions 1.3.5 and 1.3.8],  [Kempf, Theorem 1, Chapter 1, Section 1].
The affine toric variety defined by a pair $(M,\sigma)$, where $M$ is a lattice in $\mathbb{Q}^{n}$ and $\sigma\subset\mathbb{R}^{n}$ is normal.
Definition 2.5.
Let $S$ be a pointed affine semigroup. The maximal ideal of $\mathbb{K}[S]$ generated by $\{x^{\alpha}\ |\ \alpha\in S\setminus\{0\}\}$ is said to be the special ideal. The point of the toric variety that corresponds to the special maximal ideal is also called special.
3. Analytical Equivalence
Definition 3.1.
Let $\mathbf{m}$ be a maximal ideal of a ring $\mathcal{R}$. The completion of $\mathcal{R}$ with respect to $\mathbf{m}$ is the inverse limit
[…]
Let $S$ be a pointed monoid. We denote by $\mathbb{K}[[S]]$ the ring
$\mathbb{K}[[S]]:=\Big\{\sum_{\alpha\in\Lambda}a_{\alpha}X^{\alpha}\ \Big|\ a_{\alpha}\in\mathbb{K},\ \Lambda\subseteq S\Big\}.$
Note that if $S$ is not pointed, multiplication in $\mathbb{K}[[S]]$ is not defined.
Remark 3.2.
The completion of $\mathbb{K}[S]$ with respect to its maximal ideal $\mathfrak{m}=\langle X^{\alpha}\mid\alpha\in S,\alpha\neq 0\rangle$ is $\mathbb{K}[[S]]$, see [eisenbud2013commutative, Example 10.14, p. 181]. Since excellent rings satisfy that normalization and completion commute, see [egaiv1964grothendieck, Proposition 7.8.3.(vii)], it follows from Lemma 2.4 that
$\mathcal{N}(\mathbb{K}[[S]])=\mathbb{K}[[\bar{S}]].$
Lemma 3.3.
Let $\gamma^{(1)},\ldots,\gamma^{(g)}\in(\mathbb{Q}_{\geq 0})^{n}$ and let $S$ be the affine semigroup generated by $(\mathbb{Z}_{\geq 0})^{n}$ together with $\gamma^{(1)},\ldots,\gamma^{(g)}$. Then
$\mathbb{K}[[X]][X^{\gamma^{(1)}},\ldots,X^{\gamma^{(g)}}]=\mathbb{K}[[S]].$
Proof.
We proceed by induction on $g$. For $g=1$, let $S$ be generated by $(\mathbb{Z}_{\geq 0})^{n}$ and $\gamma=\beta/k$ with $\beta\in(\mathbb{Z}_{\geq 0})^{n}$ and $k>0$.
Clearly,
$\mathbb{K}[[X]][X^{\gamma}]\subseteq\mathbb{K}[[S]].$
Conversely, let $f\in\mathbb{K}[[S]]$. Then $f$ can be written as
$f=\sum_{\begin{subarray}{c}\alpha\in(\mathbb{Z}_{\geq 0})^{n}\\ m\in\mathbb{Z}_{\geq 0}\end{subarray}}a_{\alpha,m}X^{\alpha}(X^{\gamma})^{m}.$
Write $m=j+k\ell$ with $0\leq j<k$ and $\ell\in\mathbb{Z}_{\geq 0}$. Then
$f=\sum_{j=0}^{k-1}X^{j\gamma}\sum_{\ell\geq 0}\sum_{\alpha}a_{\alpha,j+k\ell}X^{\alpha}(X^{\gamma})^{k\ell}.$
Since $(X^{\gamma})^{k\ell}=X^{\beta\ell}\in\mathbb{K}[[X]]$, it follows that each inner sum lies in $\mathbb{K}[[X]]$, so $f\in\mathbb{K}[[X]][X^{\gamma}]$.
The general case follows by induction on $g$. ∎
Example 3.4.
We have
$\mathbb{K}[[x]][x^{1/2}]=\mathbb{K}[[x^{1/2}]].$
Indeed, any $y\in\mathbb{K}[[x^{1/2}]]$ can be written as
$y=\sum_{i=0}^{\infty}a_{i}(x^{1/2})^{i}=\sum_{i\text{ even}}a_{i}x^{i/2}+\sum_{i\text{ odd}}a_{i}x^{i/2}=\sum_{i\text{ even}}a_{i}x^{i/2}+x^{1/2}\sum_{i\text{ odd}}a_{i}x^{(i-1)/2},$
which lies in $\mathbb{K}[[x]][x^{1/2}]$.
Two singularities $p\in\mathbb{X}$ and $q\in\mathbb{Y}$ are said to be analytically isomorphic if there exists an isomorphism of $\mathbb{K}$-algebras
$\hat{\mathcal{O}}_{p}\simeq\hat{\mathcal{O}}_{q}.$
For example (Hartshorne, Example 5.6.3 [hartshorne2013algebraic]), let $\mathbb{X}=V(x^{2}(x+1)-y^{2})$ and $\mathbb{Y}=V(xy)$. Then the singularity $(0,0)$ in $\mathbb{X}$ is analytically isomorphic to $(0,0)$ in $\mathbb{Y}$. Geometrically, near the origin both are the union of two nonsingular branches with distinct tangents.
4. Puiseux Hypersurfaces
Definition 4.1.
A monic polynomial $f\in\mathbb{K}[[X]][y]$ is called Puiseux polynomial when there exist elements $\xi_{1},\ldots,\xi_{d}$ of the ring of Puiseux series
$\mathbb{K}[[X^{*}]]:=\bigcup_{m\in\mathbb{Z}_{>0}}\mathbb{K}\left[\left[x_{1}^{1/m},\ldots,x_{n}^{1/m}\right]\right],$
such that $f=\Pi_{i=1}^{d}(y-\xi_{i})$.
Lemma 4.2.
Let $f\in\mathbb{K}[[X]][y]$ be an irreducible Puiseux polynomial and let $\xi$ be a root of $f$. There exists an isomorphism between $\mathbb{K}[[X]][y]/\langle f\rangle$ and $\mathbb{K}[[X]][\xi]$.
Proof.
Let us consider the morphism
$\begin{split}\psi:\ \mathbb{K}[[X]][y]&\to\mathbb{K}\ [[X]][\xi(x)]\\ h(y)&\mapsto h(\xi).\end{split}$
We have $\langle f\rangle=\text{Ker}(\psi)$. Indeed, it is clear that $\langle f\rangle\subseteq\text{Ker}(\psi)$, now take $h\in\text{Ker}(\psi)$, the division algorithm implies that there exist $q,r\in\mathbb{K}[[X]][y]$ with $\text{deg}_{y}(r)<\text{deg}_{y}(f)$ such that $h=qf+r$. Since $\xi$ is a common root of $f$ and $h$, it is also a root of $r$. Since $f$ is the minimal polynomial of $\xi$, we obtain that $r$ is identically zero, and thus $\text{Ker}(\psi)\subseteq\langle f\rangle$. Together with the fact that $\psi$ is surjective, we obtain
$\mathbb{K}[[X]][y]/\langle f\rangle\simeq\mathbb{K}[[X]][\xi].$
Definition 4.3.
Let $\xi\in\mathbb{K}[[X^{*}]]$. $E=\left\{\lambda^{(1)},\ldots,\lambda^{(g)}\right\}\subset\text{Supp}\left(\xi\right)$ is a set of distinguished [^1] Given a Puiseux series $\xi\in\mathbb{K}[[X^{*}]]$, in [tornero2008kummer] the elements of a set $\left\{k\lambda^{(1)},\ldots,k\lambda^{(g)}\right\}\subset k\ \text{Supp}\left(\xi\right)\subset(\mathbb{Z}_{\geq 0})^{n}$ such that $\lambda^{(1)},\ldots,\lambda^{(g)}$ satisfy the equation (2) are called distinguished exponents. exponents of $\xi$ if
(2) $\mathbb{K}((X))\left(\xi\right)=\mathbb{K}((X))\left(X^{\lambda^{(1)}},\ldots,X^{\lambda^{(g)}}\right).$
In [tornero2008kummer], Tornero proves that all Puiseux series $\xi$, with minimal polynomial $f$ such that char $(\mathbb{K})$ does not divide  $\text{ deg }(f)$, has a set of distinguished exponents.
Given a vector with rationally independent coordinates $\omega\in\big(\mathbb{R}_{\geq 0}\big)^{n}$, let us consider the total order $\leq_{\omega}$ in $\big(\mathbb{Q}_{\geq 0}\big)^{n}$ defined by
$\lambda\leq_{\omega}\beta\hskip 14.45377pt\text{ if and only if }\hskip 14.45377pt\lambda\cdot\omega\leq\beta\cdot\omega.$
A set $\left\{\lambda^{(1)},\ldots,\lambda^{(g)}\right\}$ of distinguished exponents of a Puiseux power series $\xi$ can be constructed as follows:
(3) $\begin{split}\lambda^{(1)}&=\min_{\leq_{\omega}}\left\{\text{Supp}(\xi)\setminus\mathbb{Z}^{n}\right\}\\ \lambda^{(2)}&=\min_{\leq_{\omega}}\{\text{Supp}(\xi)\setminus(\mathbb{Z}^{n}+\lambda^{(1)}\mathbb{Z})\}\\ \vdots\\ \lambda^{(i)}&=\min_{\leq_{\omega}}\left\{\text{Supp}(\xi)\setminus(\mathbb{Z}^{n}+\sum_{j=1}^{i-1}\lambda^{(j)}\mathbb{Z})\right\}\end{split}$
for $i=2,\ldots,g$ with $g=\min\{i\ |\ \lambda^{(i)}\neq\infty\}.$ For more details we refer to [tornero2008kummer].
[…]
5. Main Results
Lemma 5.1.
Let $\lambda^{(1)},\ldots,\lambda^{(g)}\in\big(\frac{1}{k}\mathbb{Z}_{\geq 0}\big)^{n}$ be a set of distinguished exponents of a root $\xi$ of a Puiseux polynomial. There exists a $\mathbb{K}$-algebra isomorphism between the normalization of $\mathbb{K}[[X]][\xi]$ and $\mathbb{K}[[\bar{S}]]$, where $\bar{S}$ is the saturation of the affine semigroup $S$ generated by $\big(\mathbb{Z}_{\geq 0}\big)^{n}$ and $\lambda^{(1)},\ldots,\lambda^{(g)}$.
Proof.
Lemma 3.3 states that $\mathbb{K}[[X]][X^{\lambda^{(1)}},\ldots,X^{\lambda^{(g)}}]=\mathbb{K}[[S]]$, consequently
$\mathcal{N}(\mathbb{K}[[X]][X^{\lambda^{(1)}},\ldots,X^{\lambda^{(g)}}])=\mathcal{N}(\mathbb{K}[[S]])=\mathbb{K}[[\bar{S}]],$
where the second equality follows from Remark 3.2.
For each distinguished exponent $\lambda^{(i)}$, the monomial $X^{\lambda^{(i)}}$ is the root of the monic polynomial
$y^{k}-X^{k\lambda^{(i)}}\in\mathbb{K}[[X]][y]\subseteq\mathbb{K}[[X]][\xi][y].$
By definition of distinguished exponents (see equation (2)), $\mathbb{K}[[X]][\xi]$ and $\mathbb{K}[[S]]$ have the same field of fractions. Thus $X^{\lambda^{(i)}}\in\text{Frac }(\mathbb{K}[[X]][\xi])$, which implies $X^{\lambda^{(i)}}\in\mathcal{N}(\mathbb{K}[[X]][\xi])$; consequently, $\mathbb{K}[[\bar{S}]]\subseteq\mathcal{N}(\mathbb{K}[[X]][\xi])$.
On the other hand, $\xi\in\text{Frac }(\mathbb{K}[[S]])$ is a root of a monic polynomial with coefficients in $\mathbb{K}[X]\subseteq\mathbb{K}[[S]]$, hence $\xi\in\mathbb{K}[[\bar{S}]]$, completing the proof. ∎
We say that a cone is simplicial when it has a linearly independent system of generators. A singular point of a normal irreducible algebroid hypersurface in $\mathbb{A}_{\mathbb{K}}^{n+1}$ is called a Hirzebruch-Jung singularity when it is analytically isomorphic to the special point of an affine toric variety defined by a pair $(M,\sigma)$, where $M\supseteq\mathbb{Z}^{n}$ is a lattice in $\mathbb{Q}^{n}$ and $\sigma\subset\mathbb{R}^{n}$ is a simplicial cone of maximal dimension.
Theorem 5.2.
The normalization of an irreducible Puiseux hypersurface is either non-singular or has a Hirzebruch-Jung singularity. Conversely, every Hirzebruch-Jung singularity is analytically isomorphic to the normalization of a Puiseux hypersurface.
Proof.
Let $f\in\mathbb{K}[[X]][y]$ be an irreducible monic Puiseux polynomial and $\xi$ a root of $f$. Lemma 4.2 gives
$\mathbb{K}[[X]][y]/\langle f\rangle\simeq\mathbb{K}[[X]][\xi].$
Choose a set $E$ of distinguished exponents of $\xi$ as in (3) and let $S$ be the affine semigroup generated by $E$ and $(\mathbb{Z}_{\geq 0})^{n}$. Lemma 1.4 states that the saturation $\bar{S}$ is the intersection of $\mathbb{Z}S$ and the simplicial cone $(\mathbb{R}_{\geq 0})^{n}$. By Proposition 2.1, $\mathbb{K}[\bar{S}]$ is the coordinate ring of an affine toric variety, which is normal by Lemma 2.4 and its completion is $\mathbb{K}[[\bar{S}]]$ by Remark 3.2. Lemma 5.1 implies that $\mathcal{N}(\mathbb{K}[[X]][\xi])=\mathbb{K}[[\bar{S}]]$. Hence, $\mathcal{N}(\mathbb{K}[[X]][y]/\langle f\rangle)$ is either non-singular or has a Hirzebruch-Jung singularity.
Conversely, let a Hirzebruch-Jung singularity be of the form $\operatorname{Spec}(\mathbb{K}[S])$, where $S=L\cap(\mathbb{R}_{\geq 0})^{n}$ is an affine semigroup and $L$ a subgroup of $\mathbb{Z}^{n}$. Let $m_{i}\in\mathbb{Z}_{\geq 0}$ be defined by
$m_{i}e_{i}=\min\{L\cap\langle e_{i}\rangle\setminus\{0\}\},$
where $e_{i}$ is the $i$-th canonical basis vector. Define
$L^{\prime}=\left\{\left(\frac{\alpha_{1}}{m_{1}},\ldots,\frac{\alpha_{n}}{m_{n}}\right)\ \middle|\ \alpha\in L\right\},\quad S^{\prime}=L^{\prime}\cap(\mathbb{R}_{\geq 0})^{n}.$
[…]
(4) $\mathbb{K}[S^{\prime}]\simeq\mathbb{K}[S].$
Choose $\omega\in(\mathbb{R}_{\geq 0})^{n}$ with rationally independent coordinates and construct
$\lambda^{(1)}=\min_{\leq_{\omega}}\{S^{\prime}\setminus\mathbb{Z}^{n}\},\quad\lambda^{(i)}=\min_{\leq_{\omega}}\left\{S^{\prime}\setminus\mathbb{Z}^{n}+\sum_{j=1}^{i-1}\mathbb{Z}\lambda^{(j)}\right\},\quad i=2,...,g,$
with $g=\min\{i\mid\lambda^{(i)}\neq\infty\}$. Then $S^{\prime}$ is the saturation of the affine semigroup generated by $E=\{\lambda^{(1)},\ldots,\lambda^{(g)}\}$ and $(\mathbb{Z}_{\geq 0})^{n}$. Let $\xi=\sum_{i=1}^{g}X^{\lambda^{(i)}}$ and $f$ its minimal polynomial. Then
$\mathcal{N}(\mathbb{K}[[X]][y]/\langle f\rangle)\simeq\mathcal{N}(\mathbb{K}[[X]][\xi])\simeq\mathbb{K}[[S^{\prime}]]\simeq\mathbb{K}[[S]],$
where the first isomorphism is given by Lemma 4.2, the second isomorphism follows from Lema  5.1, and the third one is by (4). That proves that the given Hirzebruch-Jung singularity is analytically isomorphic to the Puiseux hypersurface defined by $f$. ∎
Example 5.3.
Lipman proved in the complex case that every normal quasi-ordinary singularity is isomorphic to a germ of the form $y^{k}-x_{1}\cdots x_{n}$ for $k\geq 2$, see [lipman1988topological, Remark 7.3.2]. In other words, every normal quasi-ordinary singularity is analytically isomorphic to the affine toric variety $\mathrm{Spec}(\mathbb{C}[S])$, where $S$ is the affine semigroup generated by $(1/k,\ldots,1/k)$ and $(\mathbb{Z}_{\geq 0})^{n}$.
Example 5.4.
Let us consider the quasi-ordinary singularity
$f(x_{1},x_{2},x_{3},y)=y_{4}-2(x_{1}^{3}x_{2}^{2}+x_{1}^{4}x_{2}^{3}x_{3}^{2})y^{2}+x_{1}^{8}x_{2}^{6}x_{3}^{4}-2x_{1}^{7}x_{2}^{5}x_{3}^{2}+x_{1}^{6}x_{2}^{4}.$
Its characteristic exponents are $(3/2,1,0)$ and $(2,3/2,1)$, see [Ayala2025]. The saturation of the affine semigroup $S$ generated by these exponents and $(\mathbb{Z}_{\geq 0})^{3}$ is $\frac{1}{2}\mathbb{Z}_{\geq 0}\times\frac{1}{2}\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}\simeq(\mathbb{Z}_{\geq 0})^{3}$. Therefore
$\mathcal{N}\left(\frac{\mathbb{K}[x_{1},x_{2},x_{3}][y]}{\langle f\rangle}\right)\simeq\mathcal{N}(\mathbb{K}[S])\simeq\mathbb{K}[x_{1},x_{2},x_{3}].$
Example 5.5.
The roots of $f=(z^{2}-x-y)^{2}-4xy\in\mathbb{K}[[x,y]][z]$ are the Puiseux series $\pm x^{1/2}\pm y^{1/2}$, which are not quasi-ordinary, see [Ayala2025]. Following Tornero’s algorithm, we obtain distinguished exponents $(1/2,0)$ and $(0,1/2)$. The affine semigroup $S$ generated by $(\mathbb{Z}_{\geq 0})^{2}$ and these exponents is already saturated. The normalization of the hypersurface defined by $f$ is analytically isomorphic to the non-singular toric surface $\mathbb{K}[S]\simeq\mathbb{K}^{2}$.
In the complex-analytic case, Popescu-Pampu [popescu2003higher, Proposition 3.5] proved that Hirzebruch-Jung singularities are precisely the normal quasi-ordinary singularities. Hence, we have:
Corollary 5.6.
The germs of normal quasi-ordinary singularities are precisely the germs of the normalizations of complex analytic irreducible Puiseux hypersurfaces.
Example 5.7.
Let $\xi=x+x^{2/3}y^{1/3}+y$ and $f$ be its minimal polynomial. The only distinguished exponent of $\xi$ is $(2/3,1/3)$. The saturation of the affine semigroup $S$ generated by $(2/3,1/3)$ and $(\mathbb{Z}_{\geq 0})^{2}$ is generated by $\{(1,0),(1/3,2/3),(2/3,1/3),(0,1)\}$. By Theorem 5.2,
$\mathcal{N}(\mathbb{C}[[x]][y]/\langle f\rangle)$
is analytically isomorphic to the toric variety $\mathrm{Spec}(\mathbb{C}[\bar{S}])$, whose canonical embedding in $\mathbb{C}^{4}$ is the Veronese variety
$\mathrm{Spec}(\mathbb{C}[\bar{S}])=V(xw-yz,y^{2}-xz,z^{2}-yw).$
The projection $\pi:(x,y,z,w)\mapsto(x,w)$ is a quasi-ordinary projection. Therefore, starting from a Puiseux hypersurface, we obtain its normalization, which is a non-hypersurface quasi-ordinary singularity.
Declarations
Competing Interests: The authors have no relevant financial or non-financial interests to disclose.
Affiliations